OpenStudy (anonymous):

METAL AND FAN! NEED HELP! Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. A) Y=-x+6 B) Y= x-6 C) Y= -x-2 D) Y= x+2 I THINK IT IS A! AM I correct???

OpenStudy (anonymous):

@iGreen @andrewhaze @cxxonicole98 @MrNood @Sunnyo Please help

OpenStudy (anonymous):

@TheEdwardsFamily @wasifa04 @HazelLuv99 @TeenWolfGirl @triciaal Please help :)

OpenStudy (theedwardsfamily):

idk

OpenStudy (anonymous):

But your a HUMANCALCULATOR! ;O @TheEdwardsFamily

OpenStudy (hazelluv99):

LOL IDK EITHER

OpenStudy (triciaal):

the slope of the given line is -1

OpenStudy (anonymous):

Yes

OpenStudy (triciaal):

perpendicular lines the product of the slope is -1 m1*m2 = -1 where m1 and m2 are the slopes the line perpendicular to this will have a slope of -1/-1 = 1

OpenStudy (anonymous):

Got it so far...

OpenStudy (triciaal):

the line passes through the point (-4, -2) y = mx + b -2 = 1(-4) + b b = 2 for any point (x, y) y = x + 2

OpenStudy (anonymous):

\[y=-x+6\]\[\implies m_{1}=-1\]Slope of the line which is perpendicular to a line is given by negative reciprocal of the slope of the line to which the line is perpendicular to \[m_{2}=-\frac{1}{-1}=1\] Now you are given the slope(1) and a point(-4,-2) through which the line is passing, using the point slope form to find the equation \[y-y_{o}=m(x-x_{o})\]

OpenStudy (anonymous):

OOOhhh I see! Thank you =) So the answer would be D

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

And @Nishant_Garg Point slope fourmula is \[y-y1=m(x-x1)\] Are those little circles a substitution for 1? xD

OpenStudy (anonymous):

It's the same thing, just how you present it, you can use any variable